OpenStudy (spinex12):

help fast! How much simple interest will be due on a loan of $9000 at a 4.25% annual interest rate for one year? $9382.50 $42.50 $3825 $382.50

OpenStudy (spinex12):

Find the simple interest. principal = $750; rate = 14%; time = 4.5 years $1222.50 $472.50 $420 $105

OpenStudy (spinex12):

Based on simple interest, find the total amount saved or owed. principal = $3000; rate = 12%; time = 8 years $3360 $5000 $5880 $2880

OpenStudy (spinex12):

The simple interest on an investment of $8000 for one year is $360. What is the interest rate? 45% 0.45% 4.5% 0.045%

OpenStudy (spinex12):

Sarah paid $30 in interest to borrow $4800 for 3 months. What annual interest rate did Sarah pay? 0.625% 1.875% 7.5% 2.5%

OpenStudy (tkhunny):

What are you doing? Would you like us to do your exam for you? Show your work on ONE problem and someone will be happy to discuss it with you.

OpenStudy (spinex12):

plz help I don't know any of this

OpenStudy (pawanyadav):

Well the Simple Interest is= P*R*T/100 Where P=principal amount R=rate T=Time In every question you have to just plug the values.

OpenStudy (pawanyadav):

Well the Simple Interest is= P*R*T/100 Where P=principal amount R=rate T=Time In every question you have to just plug the values.

OpenStudy (tkhunny):

They are all essentially the same. You need to know one thing. I = Prt I = Interest for the specified period P = Principal deposit at the beginning of the period. r = Annual Simple Interest Rate t = Time - The number of years the deposit sat in the account. 1) Write out that formula. 2) Fill in the values you know from the problem statement. 3) Solve for whatever is left. Go! Show the first one.